OpenStudy (anonymous):

would (3, 11) and (−3, −7) be the answer to What are the solutions to the following system of equations? y = x2 + 3x − 7 3x − y = −2

OpenStudy (whpalmer4):

Here's how you find out: plug \((x=3,y=11)\) into both equations. Do they both work with those values? If so, it is a solution.

OpenStudy (anonymous):

i think so

OpenStudy (whpalmer4):

Here's what I mean: \[y = x^2+3x-7\]Plug in \(x-3,y=11\) \[11=(3)^2+3(3) - 7\]\[11=9+9-7\]\[11=11\checkmark\]\[3x-y=-2\]\[3(3)-11=-2\]\[9-11=-2\]\[-2=-2\checkmark\] That shows the point \((3,11)\) is a solution to the system.

OpenStudy (whpalmer4):

The other one is correct as well. Good job!

OpenStudy (whpalmer4):

Here's a graph showing the line and the parabola and their intersection at the solutions.